The sum of two number is 112. The smaller is 58 less than the greater. Find the numbers
Found 2 solutions by Alan3354, fractalier:
Answer by Alan3354(69443)

The sum of two number is 112. The smaller is 58 less than the greater. Find the numbers
-------------
x + y = 112
x - y = 58
------------------ Add

Answer by fractalier(6550)

Okay...call the two numbers, larger and smaller, x and y. Then we have
x + y = 112 and
y = x - 58 or x - y = 58
Then we can add these together and get
x + y = 112
x - y = 58
----------------
2x = 170
x = 85 and then
y = 85 - 58 = 27
The numbers are 85 and 27.